# A primer on quasi-convex functions in nonlinear potential theories

### v2

@@ 3.6. An area formula proof of the Jensen–Słodkowski theorem
Notice that, without loss of generality, we can assume that $\delta$ defined by (3.9) is positive because if if $\delta\leqslant 0$ then $|B_{\delta}|=0$ and the inequality (3.7) is trivial. This also underlines the meaning of the estimate (3.7); one has an upper bound on the value of positive interior maximum value for $u$.
Exploting the argument above, it suffices to show that $B_{\delta}\subset Du(E)$. We will, in fact, show a weaker inclusion, but one which is still sufficient to use the argument. More precisely, consider the a priori enlargement of $E$ defined by (3.8) by eliminating the request that $u$ be twice differentiable; that is,
$\widetilde{E}:=\mathcal{C}^{+}(u,\Omega,0).$ (3.11)
The gradient map $Du$ is well defined on $\widetilde{E}$ since $\mathcal{C}^{+}(u,\Omega,0)\subset\operatorname{Diff}^{1}(u)$ by property (D at UCP) of the Upper Contact Jet Theorem 2.12. With $u,\delta,E$ and $\tilde{E}$ as above, in the following two lemmas, we will show that
@@
Let $u,E$ and $\widetilde{E}$ be as above. Then $Du(\tilde{E})\setminus Du(E)$ is null; that is, it has zero Lebesgue measure.
@@
We will prove that the larger set $Du(\widetilde{E}\setminus E)$ is null. First, note that $u$ being locally quasi-convex implies that $\widetilde{E}\setminus E$ is null by Alexandrov’s Theorem 2.2. This suggests using the well-known fact that Lipschitz maps on ${\mathbb{R}}^{n}$ send null sets to null sets. To this end, the local quasi-convexity of $u$ also implies that there is an exhaustion of $\Omega$ by compact convex sets $\{K_{i}\}_{i\in\mathbb{N}}$ such that $u$ is $\lambda_{i}$-quasi-convex on $K_{i}$. For each $i\in\mathbb{N}$, set
$E^{{(i)}}\vcentcolon=\widetilde{E}\cap K_{i}\subset\mathcal{C}^{+}(u,K_{i},0)\quad\text{and}\quad u_{i}\vcentcolon=u+\frac{\lambda_{i}}{2}|\cdot|^{2},$
which is convex on $K_{i}$.
Since $\mathcal{C}^{+}(u,K_{i},0)=\mathcal{C}^{+}(u_{i},K_{i},\lambda_{i}I)$, we know that the upper vertex map $V_{1/\lambda_{i}}$ for $u_{i}$ is well-defined and $1$-Lipschitz on $E^{{(i)}}$ (by Remark 3.17 and Proposition 3.18); that is, with $V_{1/\lambda_{i}}:=I-\lambda_{i}^{-1}Du_{i}$ one has
$\left|\left(I-\frac{1}{\lambda_{i}}Du_{i}\right)(p)-\left(I-\frac{1}{\lambda_{i}}Du_{i}\right)(q)\right|\leqslant|p-q|,\ \ \forall p,q\in{\mathbb{R}}^{n}.$
This implies that $Du_{i}$ is $2\lambda_{i}$-Lipschitz, and thus $Du$ is $3\lambda_{i}$-Lipschitz on $E^{{(i)}}$. We can now conclude that $Du(E^{{(i)}}\setminus E)$ is null for every $i\in\mathbb{N}$, by using the fact noted above that Lipschitz functions map null sets into null sets (see, e.g., [18, Theorem 2.8(i)]). The claim now follows from the countable subadditivity of the Lebesgue measure. ∎
@@
For each $p\in B_{\delta}$, set $k_{p}:=\max_{\overline{\Omega}}\left(u(\cdot)-\langle p,\cdot\rangle\right)$ which is finite since $u(\cdot)-\langle p,\cdot\rangle\in\operatorname{USC}(\overline{\Omega})$. Let $x_{p}\in\overline{\Omega}$ be a point realizes this maximim and hence one has
$u(y)\leqslant u(x_{p})+\langle p,y-x_{p}\rangle,\quad\forall\,y\in\overline{\Omega}.$ (3.13)
@@
then (3.13) says that $x_{p}$ a contact point of type 00 for $u$ on $\Omega$; that is, $u\in\widetilde{E}=\mathcal{C}^{+}(u,\Omega,0)$. Hence, by Theorem 2.12(D at UCP), $u$ is differentiable at $x_{p}$ with $Du(x_{p})=p$ so that $p\in Du(\widetilde{E})$, which proves the lemma since $p\in B_{\delta}$ was arbitrary.
@@
$\sup_{\overline{\Omega}}u=u(y)\leqslant x_{p}+\langle p,y-x_{p}\rangle\leqslant\sup_{\partial\Omega}u^{+}+|p|\,{\rm diam}(\Omega),$
and hence $|p|\geqslant\left(\sup_{\overline{\Omega}}u-\sup_{\partial\Omega}u^{+}\right)/{\rm diam}(\Omega):=\delta$, which contradicts $p\in B_{\delta}$. ∎
@@
We begin by recalling that Remark 3.29 noted that the HL–Jensen Lemma 3.23 is an equivalent formulation of the Jensen–Słodkowski Theorem 3.28. Hence, we will prove that for all $\rho>0$ small, $|\mathcal{C}^{+}(u,\overline{B}_{\rho}(x),0)|>0$ assuming that $u$ is a locally quasi-convex with strict upper contact jet $(0,0)$ at $x$. By translating $u$ we can assume that $u(x)>0$ and hence for all $\rho>0$ small, the radius $\delta$ defined by (3.9) is positive. By (3.10) with $\Omega=B_{\rho}(x)$, for all sufficiently small $\rho>0$ one has
$0<|B_{\delta}|\leqslant\int_{E}|\det D^{2}u(y)|\,\mathrm{d}y\quad\text{with}\ E:=\mathcal{C}^{+}(u,\Omega,0)\cap\operatorname{Diff}^{2}(u),$
@@
The “paraboloidal” proof of the Jensen–Słodkowski theorem passed through a density estimate for a certain set of global upper contact points for a convex function (Lemma 3.20); likewise, we can deduce a similar estimate, for the set of flat global upper contact points for a quasi-convex function on small balls around a strict local maximum point, using Alexandrov’s maximum principle.
@@
Let $u\colon X\to{\mathbb{R}}^{n}$ be $\lambda$-quasi-convex, with strict upper contact jet $(0,0)$ at $x$. Then there exists $\bar{\rho}>0$ such that
@@
Consider $y\in E:=\mathcal{C}^{+}(u,\Omega,0)\cap\operatorname{Diff}^{2}(u)$. By the quasi-convexity of $u$ we have $D^{2}u(y)\geqslant-\lambda I$, for some $\lambda>0$, and since $y$ is a maximum point of $u+\langle p,\cdot\rangle$ for some $p\in{\mathbb{R}}^{n}$, we have $D^{2}u(y)\leqslant 0$. Hence $|\det D^{2}u|\leqslant\lambda^{n}$ on $E$, and, letting $\delta$ be as in (3.9), Theorem 3.36 with $\Omega=B_{\rho}(x)$ (along with Theorem 2.2) yields $|B_{\delta}|\leqslant\lambda^{n}|\widetilde{E}|$, with $\widetilde{E}=\mathcal{C}^{+}(u,\overline{B}_{\rho}(x),0)$. The desired conclusion now is obtained by noting that
@@
4. Quasi-convex approximation of USC functions
In this section we present the well-known quasi-convex approximation of upper semicontinuous functions by way of the supconvolution. This approximation is of crucial importance for both general potential theories and the viscosity theory for PDEs since it allows one to use information on upper contact jets along the approximating sequences.
@@ 4.1. Elementary properties of semicontinuous functions
We begin by recalling the definitions and some well-known and useful properties of semicontinuous functions. Here and below $\accentset{\mbox{\large.}}{B}_{\rho}(x_{0}):=B_{\rho}(x_{0})\setminus\{x_{0}\}$ denotes the punctured open ball of radius $\rho>0$ and center $x_{0}$.
@@
    $u(x_{0})\geqslant\limsup_{X\ni x\to x_{0}}u(x):=\lim_{\rho\to 0^{+}}\sup_{x\in\accentset{\mbox{\large.}}{B}_{\rho}(x_{0})\cap X}u(x)=\inf_{\rho>0}\sup_{x\in\accentset{\mbox{\large.}}{B}_{\rho}(x_{0})\cap X}u(x)$ (4.1)
@@
    $u(x_{0})\leqslant\liminf_{X\ni x\to x_{0}}u(x):=\lim_{\rho\to 0^{+}}\sup_{x\in\accentset{\mbox{\large.}}{B}_{\rho}(x_{0})\cap X}u(x)=\sup_{\rho>0}\sup_{x\in\accentset{\mbox{\large.}}{B}_{\rho}(x_{0})\cap X}u(x).$ (4.2)
@@
Before recalling some well-known properties, a few remarks are in order. First, any function $u$ is both upper and lower semicontinuous at isolated points $x_{0}\in X$ since, by definition $\sup\emptyset=-\infty$ and $\inf\emptyset=+\infty$. Second, clearly $u$ is lower semicontinuous in $x_{0}$ if and only if $-u$ is upper semicontinuous in $x_{0}$ and hence one can (and we will) focus primarily on upper semicontinuous functions. Third, there are other equivalent definitions of upper semicontinuity in terms of the topology of upper and lower level sets of $u$; for example, $u\in\operatorname{USC}(X)$ in the sense of Definition 4.1 if and only if
@@
Fourth, semicontinuous functions which take values in ${\mathbb{R}}$ or [,+][$\infty$,$\infty$][ ∞ , + ∞ ] are also of use, but the choices made in Defintion 4.1 will be useful for the viscosity calculus of Part II. This choice is explained in the following remark.
@@
For upper semicontinuous functions allowing $\infty$- ∞ but excluding +$\infty$+ ∞ as possible values yields the following facts. Let $u:X\to[-\infty,+\infty)$ be upper semicontinuous on $X$.
@@
    $\hat{u}(x)\vcentcolon=\begin{cases}u(x)&\text{on $X$}\\ -\infty&\text{outside $X$}\,.\end{cases}$
    defines un upper semicontinuous extension of $u$ to all of ${\mathbb{R}}^{n}$.
@@
    For each $K\subset X$ compact, if $u\not\equiv-\infty$ admits a finite maximum on $K$ (see Proposition 4.3 (c)).
@@
Let $X\subset{\mathbb{R}}^{n}$.
@@
    A real-valued function on $X$ is continuous in $x_{0}\in X$ if and only if it is both upper and lower semicontinuous in $x_{0}$. In particular, $\operatorname{USC}(X)\cap\operatorname{LSC}(X)=C(X)$.
@@
    The sum of two (,+]($\infty$,$\infty$]( - ∞ , + ∞ ]-valued functions on $X$ which are both upper semicontinuous in $x_{0}\in X$ is upper semicontinuous in $x_{0}$. In particular, $\operatorname{USC}(X)+\operatorname{USC}(X)\subset\operatorname{USC}(X)$.
@@
    For every compact subset $K\subset X$, any upper semicontinuous function $K$ admits a maximum value $\max_{K}u\in[-\infty,+\infty)$ which is finite if $u\not\equiv-\infty$ on $K$.
@@
    Suppose that $u\in\operatorname{USC}({\mathbb{R}}^{n})$ is anti-coercive; that is, $u(x)\to-\infty$ as $|x|\to\infty$. Then $u$ is bounded above and the supremum $\sup_{{\mathbb{R}}^{n}}u$ is attained; that is, there exists $\bar{x}\in{\mathbb{R}}^{n}$ such that
    $\sup_{{\mathbb{R}}^{n}}u=\max_{{\mathbb{R}}^{n}}u=u(\bar{x}).$
@@
    The pointwise infimum of any family of upper semicontinuous functions on $X$ is upper semicontinuous; that is,
[…]
    where the index set $I$ has arbitrary cardinality. On the other hand, the pointwise supremum
    $\sup_{i\in I}u_{i}=\max_{i\in I}u_{i}\in\operatorname{USC}(X),\quad\text{provided that $I$ if finite}.$
@@
(a) It is enough to note that the continuity of $u$ in $x_{0}$
$\lim_{X\ni x\to x_{0}}u(x)\quad$
is clearly equivalent to the two inequalities (4.1) and (4.2).
@@
for any pair of functions $u$ and $v$ on $X$, and then one passes to the limit as $\rho\searrow 0$ if both $u$ and $v$ are upper semicontinuous in $x_{0}$.
(c) We can assume that $K$ is not a singleton $\{x_{0}\}$, since the conclusion is trivial in this case. Then, if $u\in\operatorname{USC}(K)$ is not identically $\infty$- ∞, consider a maximizing sequence $\{x_{j}\}_{j\in\mathbb{N}}\subset K$ such that
$u(x_{j})\to\sup_{K}u\quad\text{as $j\to\infty$}.$ (4.3)
Since $K$ is sequentially compact, there exists a point $x_{0}\in K$ and a subsequence $x_{j_{k}}\to x_{0}$ as $k\to+\infty$. Since $u\not\equiv-\infty$, by the choice of the codomain and by the semicontinuity of $u$, one has
$+\infty>u(x_{0})\geqslant\limsup_{k\to\infty}u(x_{j_{k}})=\lim_{k\to+\infty}u(x_{j_{k}})=\sup_{K}u,$ (4.4)
@@
(d) Let $M\vcentcolon=\sup_{{\mathbb{R}}^{n}}u$. By the anti-coercivity of $u$, there exists an open ball $B$, centered at the origin, such that $u|_{B^{\mathsf{c}}}<M$. Therefore $\sup_{{\mathbb{R}}^{n}}u=\sup_{\overline{B}}u$, and the conclusion follows from the previous point.
(e) For the first claim, if $v:=\inf_{i\in I}u$ then one has
@@
$u_{i}(x_{0})\geqslant\limsup_{X\ni x\to x_{0}}u_{i}(x)\geqslant\limsup_{X\ni x\to x_{0}}v(x)\quad\forall\,i\in I,$ (4.6)
and taking the infimum over $i\in I$ yields
$v(x_{0})\geqslant\limsup_{X\ni x\to x_{0}}v(x).$
The second claim of part (d) is left to the reader. ∎
@@
4.2. Quasi-convex approximation via supconvolution
One reason that quasi-convex functions are so interesting in general potential theories and the viscosity theory of nonlinear PDEs is the widely used fact that any upper semicontinuous $u$ function (which is bounded above) is a decreasing limit of a sequence of quasi-convex functions. The convergence of the quasi-convex approximation is pointwise and can be locally uniform if $u$ is more regular. This fact is well-known and for further details see, for example, [38, 14, 23]). The quasi-convex approximation of $u$ is given by supconvolutions of $u$ which are defined next.

### v3

@@ 3.6. An area formula proof of the Jensen–Słodkowski theorem
Notice that, without loss of generality, we can assume that $\delta$ defined by (3.9) is positive because if $\delta\leqslant 0$ then $|B_{\delta}|=0$ and the inequality (3.5) is trivial. This also underlines the meaning of the estimate (3.5); one has an upper bound on the value of positive interior maximum for $u$.
Exploiting the argument above, it suffices to show that $B_{\delta}\subset Du(E)$. We will, in fact, show a weaker inclusion, but one which is still sufficient to use the argument. More precisely, consider the a priori enlargement of $E$ defined by (3.6) by eliminating the request that $u$ be twice differentiable; that is,
$\widetilde{E}\vcentcolon=\mathcal{C}^{+}(u,\Omega,0).$
The map $Du$ is well-defined on $\widetilde{E}$ since $\mathcal{C}^{+}(u,\Omega,0)\subset\operatorname{Diff}^{1}(u)$ by property (D at UCP) of the Upper Contact Jet Theorem 2.12. With $u,\delta,E$ and $\tilde{E}$ as above, in the following two lemmas, we will show that
@@
Let $u$, $E$ and $\widetilde{E}$ be as above. Then $|Du(\tilde{E})\setminus Du(E)|=0$.
@@
We will prove that the larger set $Du(\widetilde{E}\setminus E)$ is null. First, note that $u$ being locally semiconvex implies that $\widetilde{E}\setminus E$ is null by Alexandrov’s Theorem 2.2. This suggests using the well-known fact that Lipschitz maps on ${\mathbb{R}}^{n}$ send null sets to null sets. To this end, the local semiconvexity of $u$ also implies that there is an exhaustion of $\Omega$ by compact convex sets $\{K_{i}\}_{i\in\mathbb{N}}$ such that $u$ is $\lambda_{i}$-semiconvex on $K_{i}$ for some $\lambda_{i}>0$. For each $i\in\mathbb{N}$, set
$E_{i}\vcentcolon=\widetilde{E}\cap K_{i}\subset\mathcal{C}^{+}(u,K_{i},0)\quad\text{and}\quad u_{i}\vcentcolon=u+\frac{\lambda_{i}}{2}|\cdot|^{2},$
which is convex on $K_{i}$. Since $\mathcal{C}^{+}(u,K_{i},0)=\mathcal{C}^{+}(u_{i},K_{i},\lambda_{i}I)$, we know that the upper vertex map $V_{1/\lambda_{i}}$ for $u_{i}$ is well-defined and $1$-Lipschitz on $E_{i}$ (by Remark 3.16 and Proposition 3.17); that is,
$\Bigl{|}\Bigl{(}I-\frac{1}{\lambda_{i}}Du_{i}\Bigr{)}(p)-\Bigl{(}I-\frac{1}{\lambda_{i}}Du_{i}\Bigr{)}(q)\Bigr{|}\leqslant|p-q|,\quad\forall\,p,q\in{\mathbb{R}}^{n}.$
This implies that $Du_{i}$ is $2\lambda_{i}$-Lipschitz, and thus $Du$ is $3\lambda_{i}$-Lipschitz, on $E_{i}$. We can now conclude that $Du(E_{i}\setminus E)$ is null for every $i\in\mathbb{N}$, by using the fact noted above that Lipschitz functions map null sets into null sets (see, e.g., [19, Theorem 2.8(i)]). The claim now follows from the countable subadditivity of the Lebesgue measure. ∎
@@
For each $p\in B_{\delta}$, consider $\max_{\overline{\Omega}}\left(u(\cdot)-\langle p,\cdot\rangle\right)$ which is finite since $u(\cdot)-\langle p,\cdot\rangle\in\operatorname{USC}(\overline{\Omega})$, and let $x_{p}\in\overline{\Omega}$ be a point realizes this maximum; that is
$u(y)\leqslant u(x_{p})+\langle p,y-x_{p}\rangle\quad\forall\,y\in\overline{\Omega}.$ (3.12)
@@
then (3.12) says that $x_{p}$ a contact point of type 00 for $u$ on $\Omega$; that is, $x_{p}\in\widetilde{E}=\mathcal{C}^{+}(u,\Omega,0)$. Hence, by Theorem 2.12(D at UCP), $u$ is differentiable at $x_{p}$ with $Du(x_{p})=p$ so that $p\in Du(\widetilde{E})$, which proves the lemma since $p\in B_{\delta}$ was arbitrary.
@@
$\sup_{\overline{\Omega}}u=u(y)\leqslant u(x_{p})+\langle p,y-x_{p}\rangle\leqslant\sup_{\partial\Omega}u^{+}+|p|\,{\rm diam}(\Omega),$
and hence $|p|\geqslant(\sup_{\overline{\Omega}}u-\sup_{\partial\Omega}u^{+})/{\rm diam}(\Omega)=\vcentcolon\delta$, which contradicts $p\in B_{\delta}$. ∎
@@
We recall that Remark 3.28 noted that the HL–Jensen Lemma 3.22 is an equivalent formulation of the Jensen–Słodkowski Theorem 3.27. Hence, we will prove that for all $\rho>0$ small, $|\mathcal{C}^{+}(u,\overline{B}_{\rho}(x),0)|>0$, assuming that $u$ is a locally semiconvex with strict upper contact jet $(0,0)$ at $x$. By translating $u$ we can assume that $u(x)>0$ and hence, for all $\rho>0$ small, the radius $\delta$ defined by (3.9) is positive. By (3.10) with $\Omega=B_{\rho}(x)$ (recall the definition (3.6) of $E(u)$), for all sufficiently small $\rho>0$ one has
$0<|B_{\delta}|\leqslant\int_{E}|\det D^{2}u(y)|\,\mathrm{d}y\quad\text{with}\ E\vcentcolon=\mathcal{C}^{+}(u,B_{\rho}(x),0)\cap\operatorname{Diff}^{2}(u),$
@@
The “paraboloidal” proof of the Jensen–Słodkowski theorem passed through a density estimate for a certain set of global upper contact points for a convex function (Lemma 3.19); likewise, we can deduce a similar estimate, for the set of flat global upper contact points for a semiconvex function on small balls around a strict local maximum point, using Alexandrov’s maximum principle.
@@
Let $u\colon X\to{\mathbb{R}}^{n}$ be $\lambda$-semiconvex, with strict upper contact jet $(0,0)$ at $x$. Then there exists $\bar{\rho}>0$ such that
@@
Consider $y\in E\vcentcolon=\mathcal{C}^{+}(u,B_{\rho}(x),0)\cap\operatorname{Diff}^{2}(u)$. By the semiconvexity of $u$ we have $D^{2}u(y)\geqslant-\lambda I$, for some $\lambda>0$, and since $y$ is a maximum point of $u+\langle p,\cdot\rangle$ for some $p\in{\mathbb{R}}^{n}$, we have $D^{2}u(y)\leqslant 0$. Hence $|\det D^{2}u|\leqslant\lambda^{n}$ on $E$, and, letting $\delta$ be as in (3.9), Theorem 3.35 with $\Omega=B_{\rho}(x)$ (along with Theorem 2.2) yields $|B_{\delta}|\leqslant\lambda^{n}|\widetilde{E}|$, with $\widetilde{E}=\mathcal{C}^{+}(u,\overline{B}_{\rho}(x),0)$. The desired conclusion now is obtained by noting that
@@
Chapter 4 Semiconvex approximation of upper semicontinuous functions
In this chapter we present the well-known semiconvex approximation of upper semicontinuous functions by way of the sup-convolution. This approximation is of crucial importance for both general potential theories and the viscosity theory for PDEs since it allows one to use information on upper contact jets along the approximating sequences. In addition, we will also give an alternative proof of the conventional Theorem on Sums of Crandall–Ishii–Lions which makes use of the upper contact jet technology developed up to this point.
@@ 4.1. Elementary properties of semicontinuous functions
We begin by recalling the definitions and some well-known and useful properties of semicontinuous functions. Here and below $\accentset{\mbox{\large.}}{B}_{\rho}(x_{0})\vcentcolon=B_{\rho}(x_{0})\setminus\{x_{0}\}$ denotes the punctured open ball of radius $\rho>0$ and center $x_{0}$.
@@
    $u(x_{0})\geqslant\limsup_{X\ni x\to x_{0}}u(x)\vcentcolon=\lim_{\rho\searrow 0}\sup_{x\in\accentset{\mbox{\large.}}{B}_{\rho}(x_{0})\cap X}u(x)=\inf_{\rho>0}\sup_{x\in\accentset{\mbox{\large.}}{B}_{\rho}(x_{0})\cap X}u(x)$ (4.1)
@@
    $u(x_{0})\leqslant\liminf_{X\ni x\to x_{0}}u(x)\vcentcolon=\lim_{\rho\searrow 0}\sup_{x\in\accentset{\mbox{\large.}}{B}_{\rho}(x_{0})\cap X}u(x)=\sup_{\rho>0}\sup_{x\in\accentset{\mbox{\large.}}{B}_{\rho}(x_{0})\cap X}u(x).$ (4.2)
@@
Before recalling some well-known properties, a few remarks are in order. First, any function $u$ is both upper and lower semicontinuous at isolated points $x_{0}\in X$ since, by definition $\sup\emptyset=-\infty$ and $\inf\emptyset=+\infty$. Second, clearly $u$ is lower semicontinuous at $x_{0}$ if and only if $-u$ is upper semicontinuous at $x_{0}$ and hence one can (and we will) focus primarily on upper semicontinuous functions. Third, there are other equivalent definitions of upper semicontinuity in terms of the topology of upper and lower level sets of $u$; for example, $u\in\operatorname{USC}(X)$ in the sense of Definition 4.1 if and only if
@@
Fourth, semicontinuous functions which take values in ${\mathbb{R}}$ or [,+][$\infty$,$\infty$][ - ∞ , + ∞ ] are also of use, but the choices made in Defintion 4.1 will be useful for the viscosity calculus of Part II. This choice is explained in the following remark.
@@
Given $u\colon X\to[-\infty,+\infty)$ upper semicontinuous on $X$, we have the following:
@@
    $\hat{u}(x)\vcentcolon=\begin{cases}u(x)&\text{on $X$}\\ -\infty&\text{outside $X$}\end{cases}$
    defines an upper semicontinuous extension of $u$ to all of ${\mathbb{R}}^{n}$;
@@
    for each $K\subset X$ compact, if $u\not\equiv-\infty$, then it admits a finite maximum on $K$ (see Proposition 4.3(c));
@@
Let $X\subset{\mathbb{R}}^{n}$. The following hold.
@@
    A real-valued function on $X$ is continuous at $x_{0}\in X$ if and only if it is both upper and lower semicontinuous at $x_{0}$. In particular, $\operatorname{USC}(X)\cap\operatorname{LSC}(X)=C(X)$.
@@
    The sum of two [,+)[$\infty$,$\infty$)[ - ∞ , + ∞ )-valued functions on $X$ which are both upper semicontinuous at $x_{0}\in X$ is upper semicontinuous at $x_{0}$. In particular, $\operatorname{USC}(X)+\operatorname{USC}(X)\subset\operatorname{USC}(X)$.
@@
    For every compact subset $K\subset X$, any upper semicontinuous function on $K$ admits a maximum value which is finite if $u\not\equiv-\infty$ on $K$.
@@
    Suppose that $X={\mathbb{R}}^{n}$ and $u\in\operatorname{USC}({\mathbb{R}}^{n})$ is anti-coercive; that is, $u(x)\to-\infty$ as $|x|\to\infty$. Then $u$ is bounded above and the supremum $\sup_{{\mathbb{R}}^{n}}u$ is attained; that is, there exists $\bar{x}\in{\mathbb{R}}^{n}$ such that $\sup_{{\mathbb{R}}^{n}}u=\max_{{\mathbb{R}}^{n}}u=u(\bar{x})$.
@@
    The pointwise infimum of any family of upper semicontinuous functions on $X$ is upper semicontinuous; that is, $\inf_{i\in I}u_{i}\in\operatorname{USC}(X)$ if $\{u_{i}\}_{i\in I}\subset\operatorname{USC}(X)$, where the index set $I$ has arbitrary cardinality. On the other hand, the pointwise supremum $\sup_{i\in I}u_{i}=\max_{i\in I}u_{i}\in\operatorname{USC}(X)$, provided that $I$ if finite.
@@
(a)  It is enough to note that the continuity of $u$ in $x_{0}$, written as $\lim_{X\ni x\to x_{0}}u(x)$, is clearly equivalent to the two inequalities (4.1) and (4.2).
@@
for any pair of functions $u$ and $v$ on $X$, and then one passes to the limit as $\rho\searrow 0$ if both $u$ and $v$ are upper semicontinuous at $x_{0}$.
(c)  We can assume that $K$ is not a singleton $\{x_{0}\}$, since the conclusion is trivial in this case. Then, if $u\in\operatorname{USC}(K)$ is not identically $\infty$- ∞, consider a maximizing sequence $\{x_{j}\}_{j\in\mathbb{N}}\subset K$; that is, $u(x_{j})\to\sup_{K}u$ as $j\to\infty$. Since $K$ is sequentially compact, there exists a point $x_{0}\in K$ and a subsequence $x_{j_{k}}\to x_{0}$ as $k\to+\infty$. Since $u\not\equiv-\infty$, by the choice of the codomain and by the semicontinuity of $u$, one has
${+\infty}>u(x_{0})\geqslant\limsup_{k\to\infty}u(x_{j_{k}})=\lim_{k\to+\infty}u(x_{j_{k}})=\sup_{K}u,$
@@
(d)  Let $M\vcentcolon=\sup_{{\mathbb{R}}^{n}}u$. By the anti-coercivity of $u$, there exists an open ball $B$, centered at the origin, such that ${\left.\kern-1.2ptu\vphantom{|}\right|_{B^{\mathsf{c}}}}<M$. Therefore $\sup_{{\mathbb{R}}^{n}}u=\sup_{\overline{B}}u$, and the conclusion follows from the previous point.
(e) For the first claim, if $v\vcentcolon=\inf_{i\in I}u$ then one has
@@
$u_{i}(x_{0})\geqslant\limsup_{X\ni x\to x_{0}}u_{i}(x)\geqslant\limsup_{X\ni x\to x_{0}}v(x)\quad\forall\,i\in I,$
and taking the infimum over $i\in I$ yields $v(x_{0})\geqslant\limsup_{X\ni x\to x_{0}}v(x)$. The second claim of part (d) is left to the reader. ∎
@@
4.2. Semiconvex approximation via sup-convolution
One reason why semiconvex functions are so interesting in general potential theories and the viscosity theory of nonlinear PDEs is the widely used fact that any upper semicontinuous $u$ function (which is bounded above) is a decreasing limit of a sequence of semiconvex functions. The convergence of the semiconvex approximation is pointwise and can be locally uniform if $u$ is more regular. This fact is well known and for further details see, for example, [43, 15, 26]. The semiconvex approximation of $u$ is given by sup-convolutions of $u$ which are defined next.